OpenStudy (anonymous):

Calc (Derivatives) am I doing these problems correctly? (teacher doesn't want us to simplify) (question in comments)

OpenStudy (anonymous):

\[y=\sqrt{4x^2 + x -3}\] \[y'(x) = 4x^2 +x -3) ^(1/2)\] then, using the power rule \[= (1/2)(4x^2 + x - 3)^(-1/2) (8x+1)\]

ganeshie8 (ganeshie8):

\(= (1/2)(4x^2 + x - 3)^{-1/2} (8x+1)\)

ganeshie8 (ganeshie8):

like that ?

OpenStudy (anonymous):

\[y=(7x-92)^\frac{ 1 }{ 4 } \] to \[\frac{ 1 }{ 4 }(7x-92)^\frac{ 3 }{ 4 }(7)\]

OpenStudy (anonymous):

yeah, sorry

ganeshie8 (ganeshie8):

first one is correct

ganeshie8 (ganeshie8):

for second one, check the sign on exponent...

OpenStudy (anonymous):

Ok, cool

ganeshie8 (ganeshie8):

\(\frac{ 1 }{ 4 }(7x-92)^\frac{ \color{red}{-}3 }{ 4 }(7) \)

OpenStudy (anonymous):

oh, yeah. I forgot to make it negative :)

OpenStudy (anonymous):

Can I then multiply (1/4) and 7?

ganeshie8 (ganeshie8):

absolutely !

OpenStudy (anonymous):

cool! thanks :)

OpenStudy (anonymous):

Tell me, how much harder does calculus get?

OpenStudy (anonymous):

oh, so, on the second Problem I have (-7/4)(7x-92)^(-3/4) when I move the negative exponent down, does the problem read \[f'(x)=\frac{ -7 }{ 4 }/[(7x-92)^\frac{ 3 }{ 4 }]\] or \[\frac{ -7 }{ 4(7x-92)^\frac{ 3 }{ 4 } }\] ?

OpenStudy (anonymous):

Sorry for formatting, they are both supposed to look like the second example

ganeshie8 (ganeshie8):

hey no

ganeshie8 (ganeshie8):

\(\large y=(7x-92)^\frac{ 1 }{ 4 } \) \(\large y'=\frac{1}{4}(7x-92)^{\frac{ 1 }{ 4 }-1} (7) \) \(\large y'=\frac{7}{4}(7x-92)^{\frac{ -3}{ 4 }} \) \(\large y'=\frac{7}{4(7x-92)^{\frac{ 3}{ 4 }}} \)

ganeshie8 (ganeshie8):

there will not be any negative signs... negative sign is in exponent only when we first differentiate..

OpenStudy (anonymous):

Thank you! I don't know where that negative came from lol. Thanks again!